If twice Allen's age is increased by 48, the result is three times Allen's age? How old is Allen?
Answer by oberobic(2304)

A = Allen's age
.
Twice Allen's age increased by 48: 2A+48
.
Three times Allen's age: 3A
.
3A = 2A+48
A = 48
.
Allen's age is 48.
.
Check this solution
3A = 3*48 = 144
2A+48 = 96+48 = 144
.
Done.